A tetrahedron has 4 vertices and 6 edges. Find the number of faces it has.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Using Euler’s formula for polyhedra,

V – E + F = 2

So, 4 – 6 + F = 2

F = 2 + 2

F = 4.
Faces = 4

Thanks